Two Trains left a station at the same time. One, Train A, traveled north at a certain speed and the other, Train B, traveled south at twice the speed. After 4 hours, the trains were 600 miles apart. How fast was each train traveling? Use needs to input the initial speed of train a to being the simulation.

Answer by checkley79(3056)

D+RT
600=(x+2x)4
600=3x*4
600=12x
600/12=x
x=50 mph ans. for the slower train.
2*50=100 mph ans. for the second train.
PROOF:
600=(50+100)4
600=150*4
600=600